sin(a+b)=3/5 cos(a-b)=12/13. find tan2a.

sin(a+b)=3/5
cos(a-b)=12/13.
find tan2a.

Grade:12

3 Answers

Aman Bansal
11 years ago

Dear Piyush,

Cos(A+B)=4/5= 0.8
A+B =36.87 

Sin(A-B)=5/13 = 0.3846
A-B = 22.62
by adding
2A =36.87 + 22.62 = 59.49
tan 2A = tan 59.49= 1.6969

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


TANAYRAJ SINGH CHOUHAN
11 years ago
sin(a+b) = 3/5 i.e. a+b = 37 & cos(a-b) = 12/13 i.e. a-b = 22.5 These are standard values of some trigonometric functions which are used as shortcuts So , from above equations we get a=30(approx) So tan2a = tan60 = v3 If you are satisfied by my answer please click yes below the answer If not satisfied post your reply back so that i can explain better & BEST OF LUCK!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
sheik arbaz
11 years ago

Dear friend,

sin(a+b)=3/5 cos(a-b)=12/13

sin,cos are possitive .so,they lies in Q1

 

 

 

(by pythagarus rule adj side is 4)

3 5

 

4 (since sin(a+B)=3/4)

 

 

=a+b

 

therefore tan(a+b)=3/4

 

cos(a-b)=12/13

 

 

5 13 (by pythagarus rule opp'' side is 5)

 

 

12 (since cos(a-b)=12/13)

=a-b

(therefore tan(a-b)=5/12)

tan(a+b+a-b)=tan2a

=tan(a+b)+tan(a-b)

1-tan(a+b).tan(a-b)

 

= 3/4 +5/12

1-3/4 *5/12

(by solving this you get tan2a=56/33)

I think you like this .

Please,please,please approve it...

gud bye.All the best.